Are ClO3(-) and SO3(-2) Isostructural?If yes then.why?

Grade:12th pass

3 Answers

Umakant biswal
7 years ago
@ shweta 
isostrctural species are those species which have same hybridisation in the central atom . 
and NO , these two are not isostrctural beacuse they have diffrent hybridisation 
in clo3- the hybridisation of cl as central atom = sp3 
in so3 2- the hybridisation of s as central atom = sp2 
so, as they have diffrent hybridisation , they are not isostrctural . 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..
Bejoi Mathew
7 years ago
Isostructural species are those species which have not only have similar hybridisation but also same lone pair and bond pair of electrons ie of the species compared.
This need not be true in the case of BeCl2 and XeF2.
NOW IN YOUR DOUBT OF WHETHER ClO3- and SO32- ARE ISOSTRUCTURAL.
YES THEY ARE AS THEY DO HAVE THE SAME HYBRIDISATION.
SEE,
ClO3-- ½ ( 7 + 1) = 4 (sp3)
SO32- - ½ (6+ 2) = 4 (sp3)
HOPE THIS HELPS.
umakant you have made a calculation error, please rectify.
 
Labba
4 years ago
actually umakant has given a correct answer but it is for so3 not for so3(-2). answer given by bejoy is accurate.